# TTS-VAR: A Test-Time Scaling Framework for Visual Auto-Regressive Generation

### v1

@@ 5.3 Resampling for Superior Samples
Analysis. In path searching, resampling intermediate states with high potentials is a straightforward yet effective approach for superior results with minimal consumption. However, in VAR, this may not be advantageous and can even be detrimental at certain scales. In Fig. 5 (a), we illustrate the score distinctions between solely employing Best-of-N ($N=2,4$) and concurrently applying potential (VALUE) resampling at specific scales. Although Best-of-N selection ensures comparatively high results, it is evident that resampling at early scales (e.g., scale 3) leads to a noticeable decline in the final results. Instead, resampling at later scales yields a certain degree of improvement.
$N$ Resampling Scale GenEval ImageReward HPS CLIP Aesthetic
@@
Table 3: Resampling Scale Difference. This table shows results with different resampling scales, indicating the influence of resampling frequency.
We analyze this phenomenon from the perspective of consistency between intermediate states and final images, as shown in Fig. 5 (b). We compute potential scores at each scale and select the one with the highest potential accordingly. We then assess whether the selected best intermediate state aligns with the best final image (the optimal state at scale 12), resulting in a sequence of scores between 0 and 1, termed consistency. Low consistency of early scales in this evolving curve indicates that those scores rarely accurately reflect the quality of final results. The scores become valuable from a certain late scale, like scale 6, with comparatively high consistency. This explains why resampling efficacy varies and should be applied selectively on later scales.
Resampling Scales. Based on the aforementioned observation, we further investigate whether increasing the resampling frequency at late scales is beneficial. As illustrated in Table 3, compared with raw inference, resampling greatly enhances the results. However, increasing the frequency has a negligible impact. For instance, there is a modest improvement in ImageReward and HPS for scales [6,8,10] compared to scales [6,9], but at the cost of Geneval. Considering that executing resampling incurs additional computational expenses related to image decoding and score calculation, we opt to resample only at scales 6 and 9.
@@
$N$ Potential GenEval ImageReward HPS CLIP Aesthetic
@@
Table 4: Potential Score Difference. This table shows results using different potential calculating means as scores for resampling.
Potential Scores. As mentioned in Sec. 4.3, we have developed distinct computational modes to explore those with higher potentials. Initially, we seek better options through a theoretical approach by visualizing the consistency. As exhibited, DIFF continuously yields low consistency levels and fails to predict the desired outcomes. SUM demonstrates a stable increment but with comparatively low values. In contrast, VALUE and MAX exhibit similar characteristics, maintaining relatively high scores since scale 6 and showing a steady increase.
Experiments in Table 4 with $N=2,4$ present coherent results. Among these potentials, DIFF lags in all indicators. Although SUM achieves some acceptable outcomes, the overall score remains low. As forecasted by consistency, VALUE and MAX achieve the highest scores in text-related metrics such as GenEval, ImageReward, and HPS, indicating a higher likelihood of selecting for superior final results. Considering that MAX requires score calculations at each scale and leads to an additional computational cost, we utilize VALUE as the potential score.
@@ 5.4 Clustering for Structural Diversity
Figure 7: Visualization of Generation Process. The text prompt here is "a photo of a bottle and a bicycle". The left is one final generated image. The right is the corresponding generation process and visualized DINOv2 features extracted from different scales. It demonstrates that features captured in early scales can indicate the structural information.
Analysis. In reference to Sec. 5.3, resampling is not universally applicable to all scales. Nevertheless, in VAR, the effectiveness and low cost associated with early scales present an invaluable and unmissable opportunity to search for more samples, thereby unlocking greater potential for the final outcomes. We notice that, given the same prompt, the structure of images significantly influences the scores. Moreover, unlike details that emerge later, structural information can be captured from the early scales. The right side of Fig. 7 demonstrates that the generation process follows a structure-to-detail progression, with rough outlines becoming discernible from scale 2. When using DINOv2 to extract intermediate images and visualizing them through PCA, as seen in the bottom line, these features exhibit characteristics akin to the original images. Consequently, we leverage structural information and conduct clustering-based diversity search to sample dissimilar structures, thus scaling for more possibilities, especially when resampling may not suffice.
$N$ Clustering Scale GenEval ImageReward
@@
$N$ Extractor GenEval ImageReward
@@
Clustering Scales. According to Fig. 7, features at scale 2 display a coarse structure, while those at scale 5 reveal a refined structure similar to final outcomes. Therefore, we specifically apply clustering on these scales. The results of $N=2,4$, with and without clustering, are presented in Table 6. The first lines of each block denote the results without clustering (Best-of-N), with subsequent lines showing outcomes under different clustering scales. Evidently, each clustering increases the likelihood of yielding better results, and there is an obvious growth when employing both scales.
Extractor. We tested various extractors for clustering features outlined in Table 6. Both PCA and Pool are transformations of features extracted by DINOv2 [30], as detailed in Sec. 4.2. While supervised InceptionV3 [48] features perform optimally in ImageReward, they notably underperform in GenEval. PCA delivers superior results on average and is employed. We attribute this to that the patch-level features from PCA align more closely with the observed structural traits.
@@ Appendix A Algorithm of TTS-VAR
We describe the algorithm of TTS-VAR in Alg. 1. Following the generation process of VAR [13] (Infinity [14]), TTS-VAR first predicts the residual tokens at the current scale and adds them to the accumulated feature maps. At scales that require clustering, TTS-VAR uses the extractor to gather features from $b_{i}$ intermediate images decoded from the feature maps. It then clusters the samples based on these features and selects $b_{i+1}$ ones as the next batch. At scales that require resampling, TTS-VAR employs the potential function to calculate scores for each image and samples $b_{i+1}$ indexes from the multinomial distribution for superior intermediate states.
@@
1:Scales $\mathcal{S}=\{s_{1},s_{2},...,s_{K}\}$, Descending batch sizes =$\mathcal{B}=\{b_{1},b_{2},...,b_{K}\}$, Clustering scales set $S_{c}$, Resampling scales set $S_{r}$, Generative model $\theta$, Reward model $r_{\phi}$ Extractor $F$, Potential Score function $P$, Text prompt $c$.
@@
Appendix B Detailed Main Results
$N$ Strategy GenEval ImageReward HPS CLIP Aesthetic
@@ Appendix B Detailed Main Results
Table 6: Scores over Different Strategies. This table shows results with different scaling strategies.
We present detailed results of variant curves in Table 6. As evident, TTS-VAR demonstrates clear advantages across all indicators [54, 53, 57, 58, 60] compared to the baselines. In Table 7, we list each item of the GenEval [54] metric. Generally, our method significantly improves performance on handling two objects and counting tasks. We attribute this to the importance of structural accuracy in multi-character scenes, particularly when two objects and multiple identical objects (counting) are involved. For instance, when provided with a prompt for three objects, there is a possibility that the model may incorrectly generate a layout with four objects. Once this error occurs, following the structure-to-detail generation process in VAR, it becomes challenging for subsequent scales to rectify. However, TTS-VAR facilitates structure diversity, thereby enabling the selection of a layout with the correct configuration and avoiding the irreversible wrong generation process for inferior samples.
$N$ Strategy Overall Single Obj. Two Obj. Counting Colors Position Color Attri.
@@
Table 7: GenEval Details. This table shows each item of the GenEval benchmark, "Object" is short for "Obj.", and "Attribute" is short for "Attri.".
Appendix C Performance over Computational Consumption
We here display the changing curves of GenEval, ImageReward, and HPSv2 over the increment of computation in Fig. 8, along with the increment of sample number $N$. As shown, our method TTS-VAR has higher computational efficiency and surpasses Importance Sampling and Best-of-N with less than half TFLOPs.
@@ Appendix C Performance over Computational Consumption
Appendix D Ablation Study
D.1 Pipeline Ablation
We present the ablation study of different design components in the overall pipeline in Table 8. As adaptive batch sampling alone (without integrated sample selection mechanisms) cannot directly enhance generation performance, these cases are denoted by "-". Excluding these baseline cases, both clustering-based diversity search and resampling-based potential selection demonstrate performance improvements, with statistically significant gains observed in reward and related evaluation indicators.
@@ D.1 Pipeline Ablation
$N$ Method GenEval ImageReward HPS CLIP Aesthetic
@@
D.2 Reward Models
@@
As shown in Table 9, generally, each reward model motivates an increase in the corresponding metric. For instance, with $N=4$, the Aesthetic model, the ImageReward model, and the HPS model achieve the highest scores in the associated indicators, respectively. Among different models, ImageReward promotes improvements more. Especially with $N=2$, ImageReward demonstrates a clear lead in GenEval and even defeats HPS in HPS score. We attribute this to the ability to clearly distinguish between superior and inferior samples, along with scoring that better aligns with human preferences.
$N$ Reward Model GenEval ImageReward HPS CLIP Aesthetic
@@
D.3 $\lambda$ Setting
In Table 10, we exhibit the results using different temperature $\lambda$ in the resampling process with fixed clustering operations. Intuitively, higher temperature promotes the expression of intermediate states with higher potential scores and prevents superior samples. However, excessively high temperatures can also widen the gap between intermediate states with the highest scores and those with scores that are only slightly lower. This can directly inhibit the generation of these slightly lagging intermediate states, which may ultimately become the optimal results. As shown, though there is a steady increase in ImageReward, $\lambda=10.0$ falls behind $\lambda=5.0$ with $N=4$ in GenEval.
$N$ Lambda GenEval ImageReward HPS CLIP Aesthetic

### v2

@@ 5.4 Resampling for Superior Samples
Analysis. In path searching, resampling intermediate states with high potentials is a straightforward yet effective approach for superior results with minimal consumption. However, in VAR, this may not be advantageous and can even be detrimental at certain scales. In Fig.˜5 (a), we illustrate the score distinctions between solely employing Best-of-N ($N=2,4$) and concurrently applying potential (VALUE) resampling at specific scales. Although Best-of-N selection ensures comparatively high results, it is evident that resampling at early scales (e.g., scale 3) leads to a noticeable decline in the final results. Instead, resampling at later scales yields a certain degree of improvement.
NN Resampling Scale GenEval ImageReward HPS CLIP Aesthetic
@@
Table 4: Resampling Scale Difference. This table shows results with different resampling scales.
We analyze this phenomenon from the perspective of consistency between intermediate states and final images, as shown in Fig.˜5 (b). We compute potential scores at each scale and select the one with the highest potential accordingly. We then assess whether the selected best intermediate state aligns with the best final image (the optimal state at scale 12), resulting in a sequence of scores between 0 and 1, termed consistency. Low consistency of early scales in this evolving curve indicates that those scores rarely accurately reflect the quality of final results. The scores become valuable from a certain late scale, like scale 6, with comparatively high consistency. This explains why resampling efficacy varies and should be applied selectively on later scales.
Resampling Scales. Based on the aforementioned observation, we further investigate whether increasing the resampling frequency at late scales is beneficial. As illustrated in Table˜4, compared with raw inference, resampling greatly enhances the results. However, increasing the frequency has a negligible impact. For instance, there is a modest improvement in ImageReward and HPS for scales [6,8,10] compared to scales [6,9], but at the cost of Geneval. Considering that executing resampling incurs additional computational expenses related to image decoding and score calculation, we opt to resample only at scales 6 and 9.
@@
NN Potential GenEval ImageReward HPS CLIP Aesthetic
@@
Table 5: Potential Score Difference. This table shows results with different potential scores.
Potential Scores. As mentioned in Sec.˜4.3, we have developed distinct computational modes to explore those with higher potentials. Initially, we seek better options through a theoretical approach by visualizing the consistency. As exhibited, DIFF continuously yields low consistency levels and fails to predict the desired outcomes. SUM demonstrates a stable increment but with comparatively low values. In contrast, VALUE and MAX exhibit similar characteristics, maintaining relatively high scores since scale 6 and showing a steady increase.
Experiments in Table˜5 with $N=2,4$ present coherent results. Among these potentials, DIFF lags in all indicators. Although SUM achieves some acceptable outcomes, the overall score remains low. As forecasted by consistency, VALUE and MAX achieve the highest scores in text-related metrics such as GenEval, ImageReward, and HPS, indicating a higher likelihood of selecting for superior final results. Considering that MAX requires score calculations at each scale and leads to an additional computational cost, we utilize VALUE as the potential score.
@@ 5.5 Clustering for Structural Diversity
Figure 7: Visualization of Generation Process. The text prompt is "a photo of a bottle and a bicycle". The left is the generated image. The right is the generation process and visualized DINOv2 features from scales. It demonstrates that features in early scales can indicate the structural information.
Analysis. In reference to Sec.˜5.4, resampling is not universally applicable to all scales. Nevertheless, in VAR, the effectiveness and low cost associated with early scales present an invaluable and unmissable opportunity to search for more samples, thereby unlocking greater potential for the final outcomes. We notice that, given the same prompt, the structure of images significantly influences the scores. Moreover, unlike details that emerge later, structural information can be captured from the early scales. The right side of Fig.˜7 demonstrates that the generation process follows a structure-to-detail progression, with rough outlines becoming discernible from scale 2. When using DINOv2 to extract intermediate images and visualizing them through PCA, as seen in the bottom line, these features exhibit characteristics akin to the original images. Consequently, we leverage structural information and conduct clustering-based diversity search to sample dissimilar structures, thus scaling for more possibilities, especially when resampling may not suffice.
NN Clustering Scale GenEval ImageReward
@@
NN Extractor GenEval ImageReward
@@
Clustering Scales. According to Fig.˜7, features at scale 2 display a coarse structure, while those at scale 5 reveal a refined structure similar to final outcomes. Therefore, we specifically apply clustering on these scales. The results of $N=2,4$, with and without clustering, are presented in Table˜7. The first lines of each block denote the results without clustering (Best-of-N), with subsequent lines showing outcomes under different clustering scales. Evidently, each clustering increases the likelihood of yielding better results, and there is an obvious growth when employing both scales.
Extractor. We tested various extractors for clustering features outlined in Table˜7. Both PCA and Pool are transformations of features extracted by DINOv2 [30], as detailed in Sec.˜4.2. While supervised InceptionV3 [47] features perform optimally in ImageReward, they notably underperform in GenEval. PCA delivers superior results on average and is employed. We attribute this to that the patch-level features from PCA align more closely with the observed structural traits.
@@ 6 Conclusion
Acknowledgements
This work is supported by the National Nature Science Foundation of China (No. 62402406).
@@ Appendix A Algorithm of TTS-VAR
We describe the algorithm of TTS-VAR in Alg.˜1. Following the generation process of VAR [13] (Infinity [14]), TTS-VAR first predicts the residual tokens at the current scale and adds them to the accumulated feature maps. At scales that require clustering, TTS-VAR uses the extractor to gather features from $b_{i}$ intermediate images decoded from the feature maps. It then clusters the samples based on these features and selects $b_{i+1}$ ones as the next batch. At scales that require resampling, TTS-VAR employs the potential function to calculate scores for each image and samples $b_{i+1}$ indexes from the multinomial distribution for superior intermediate states.
@@
1:Scales $\mathcal{S}=\{s_{1},s_{2},...,s_{K}\}$, Descending batch sizes $\mathcal{B}=\{b_{1},b_{2},...,b_{K}\}$, Clustering scales set $S_{c}$, Resampling scales set $S_{r}$, Generative model $\theta$, Reward model $r_{\phi}$ Extractor $F$, Potential Score function $P$, Text prompt $c$.
@@
Appendix B Hyperparameters Settings
In K-Means, we set the next batch size $b_{i+1}$ in the adaptive descending batch sizes as the number of centers, to find $b_{i+1}$ different structure categories. For PCA, we only select the first major component, as we utilize it for dimension reduction only. In all benchmarks, we use seeds from 0 to 3 for different results. For experiments requiring multi-GPU, like $N=8$, we set the seed from 100*process_index to 100*process_index+3 for each device. We use adaptive batch size [8NN,8NN,6NN,6NN,6NN,4NN,2NN,2NN,2NN,1NN,1NN,1NN,1NN] for Infinity2B with 1024 pixel and [8NN,8NN,6NN,6NN,6NN,4NN,2NN,2NN,1NN,1NN] for Infinity8B with 512 pixel.
For other parameters related to Infinity, we keep them the same as the original public settings.
Appendix C Detailed Main Results
NN Strategy GenEval ImageReward HPS CLIP Aesthetic
@@ Appendix C Detailed Main Results
Table 6: Scores over Different Strategies of Infinity2B.
We present detailed Infinity2B results of variant curves in Table˜6. As evident, TTS-VAR demonstrates clear advantages across all indicators [53, 52, 56, 57, 59] compared to the baselines. In Table˜7, we list each item of the GenEval [53] metric. Generally, our method significantly improves performance on handling two objects and counting tasks. We attribute this to the importance of structural accuracy in multi-character scenes, particularly when two objects and multiple identical objects (counting) are involved. For instance, when provided with a prompt for three objects, there is a possibility that the model may incorrectly generate a layout with four objects. Once this error occurs, following the structure-to-detail generation process in VAR, it becomes challenging for subsequent scales to rectify. However, TTS-VAR facilitates structure diversity, thereby enabling the selection of a layout with the correct configuration and avoiding the irreversible wrong generation process for inferior samples.
NN Strategy Overall Single Obj. Two Obj. Counting Colors Position Color Attri.
@@
Table 7: GenEval Details. This table shows each item of the GenEval benchmark. "Object" is short for "Obj.", and "Attribute" is short for "Attri.".
In Table˜8, we present the detailed different scores of Infinity8B. As shown, our method promises a stable increase in different models, including ones with superior performance.
NN Strategy GenEval ImageReward HPS CLIP Aesthetic
1 Raw Inference 0.7646 1.2095 0.3081 0.3388 0.5721
1 Ours 0.7931 1.3929 0.3115 0.3414 0.5727
2 Importance Sampling 0.7834 1.3091 0.3097 0.3399 0.5716
2 Best-of-N 0.7880 1.3435 0.3127 0.3423 0.5722
2 Ours 0.7985 1.4572 0.3106 0.3411 0.5730
4 Importance Sampling 0.7901 1.3786 0.3110 0.3409 0.5722
4 Best-of-N 0.7995 1.4092 0.3122 0.3397 0.5730
4 Ours 0.8189 1.5100 0.3139 0.3425 0.5727
Table 8: Scores over Different Strategies of Infinity8B.
Appendix D Results on DPG-Bench
We evaluate the performance of the baseline Infinity model, the Best-of-N (BoN) method, and our approach on the DPG-Bench, specifically for the case where $N=4$. For each prompt, we generated four corresponding images for evaluation using seeds 0 through 3. The results are presented in Table 9.
Methods Global Relation Overall
SDv2.1 77.67 80.72 68.09
DALL-E 3 90.97 90.58 83.50
SDXL 83.27 86.76 74.65
PixArt-Sigma 86.89 86.59 80.54
SD3 (d=24) - - 84.08
HART - - 80.89
Show-o - - 67.48
Emu3 - - 81.60
Infinity 80.96 89.83 81.72
Infinity2B+BoN ($N=4$) 88.02 87.33 82.52
Infinity2B+Ours ($N=4$) 83.37 88.81 82.94
Table 9: Evaluation results on DPG-Bench with $N=4$.
Appendix E Performance over Computational Consumption
We here display the changing curves of GenEval, ImageReward, and HPSv2 over the increment of computation in Fig.˜8, along with the increment of sample number $N$. As shown, our method TTS-VAR has higher computational efficiency and surpasses Importance Sampling and Best-of-N with less than half TFLOPs.
@@ Appendix E Performance over Computational Consumption
Appendix F Evaluation on Image Diversity
We follow [24] to adopt an evaluation on the CLIP latent space, which is more in accordance with the T2I task. The metric CLIP-Div, as formulated below, with CLIP encoder $f_{\theta}$ and $k$ samples $\{\mathbf{x}_{0}^{i}\}_{i=1}^{k}$, calculates the difference between samples with the same prompt and directly measures the diversity on text-aligned features.
$\text{CLIP-Div}\left(\{\mathbf{x}_{0}^{i}\}_{i=1}^{k}\right):=\sum_{i=1}^{k}\sum_{j=i}^{k}\frac{2}{k(k-1)}\left\|f_{\theta}(\mathbf{x}_{0}^{i})-f_{\theta}(\mathbf{x}_{0}^{j})\right\|_{2}^{2}$ (4)
Though the metric is under a relatively small range, we can judge the diversity by comparing the relative differences between different methods, as shown in Table 10. Generally, with the growing $N$, the generated results gain a higher possibility of convergence to adjacent superior features, resulting in a lower CLIP-Div score. However, the difference is quite small, indicating that our method enhances the performance with little sacrifice of diversity.
Method $N=1$ $N=2$ $N=4$ $N=8$
Importance Sampling 0.0828 0.0819 0.0822 0.0794
Best-of-N 0.0828 0.0815 0.0788 0.0791
Ours 0.0792 0.0788 0.0789 0.0779
Table 10: CLIP-Div scores under different sampling strategies and values of NN.
Appendix G Ablation Study
G.1 Pipeline Ablation
We present the ablation study of different design components in the overall pipeline in Table˜11. As adaptive batch sampling alone (without integrated sample selection mechanisms) cannot directly enhance generation performance, these cases are denoted by "-". Excluding these baseline cases, both clustering-based diversity search and resampling-based potential selection demonstrate performance improvements, with statistically significant gains observed in reward and related evaluation indicators.
@@ G.1 Pipeline Ablation
NN Method GenEval ImageReward HPS CLIP Aesthetic
@@
G.2 Computation-Performance Tradeoff
We evaluate our method with $N=2,4$ on the GPU Nvidia A800-SXM4-80GB. Below is the ablation of different designs and corresponding gains in inference time and peak memory utilization. Note that we do not include image decoding time here, as this is related to VAE and image size, which is irrelevant to the test-time scaling method itself. Because of the upper boundary of the 80G memory, BoN with $N=8$ requires two inferences and doubles the consumption.
NN Method GenEval ImageReward Inference Time (s) Peak Memory (GB)
1 Infinity 0.6946 1.1320 1.15 20.34
2 Infinity 0.6946 1.1320
+BoN 0.7087 1.2545 1.70 28.12
+Adaptive Batch Sampling - - 2.20 28.20
+Clustering 0.7220 1.2591 2.28 29.90
+Resampling (Ours) 0.7403 1.4136 2.51 30.73
4 Infinity 0.6946 1.1320
+BoN 0.7244 1.3471 2.84 43.68
+Adaptive Batch Sampling - - 3.99 43.87
+Clustering 0.7294 1.3608 4.12 45.57
+Resampling (Ours) 0.7437 1.4605 4.59 46.40
8 Infinity+BoN 0.7364 1.4144 5.70 43.68
Table 12: Computation-performance tradeoff of different designs.
Our method takes less time and memory to achieve comparable or even superior performance, for example, BoN with $N=4$ and ours with $N=2$. This proves the effectiveness and efficiency of TTS-VAR.
G.3 Reward Models
@@
As shown in Table˜13, generally, each reward model motivates an increase in the corresponding metric. For instance, with $N=4$, the Aesthetic model, the ImageReward model, and the HPS model achieve the highest scores in the associated indicators, respectively. Among different models, ImageReward promotes improvements more. Especially with $N=2$, ImageReward demonstrates a clear lead in GenEval and even defeats HPS in HPS score. We attribute this to the ability to clearly distinguish between superior and inferior samples, along with scoring that better aligns with human preferences.
NN Reward Model GenEval ImageReward HPS CLIP Aesthetic
@@
G.4 Ablation on Clustering
We compare our clustering method with a random drop process in Table 14. As shown in the results, even though resampling is in place to help ensure image quality, the effectiveness of scaling is significantly diminished when clustering is removed. This shows that clustering serves as an effective selection mechanism.
NN Method GenEval ImageReward HPS CLIP Aesthetic
1 Infinity2B+Ours 0.7253 1.3226 0.3084 0.3395 0.5822
w/o Clustering 0.7193 1.2740 0.3072 0.3391 0.5821
2 Infinity2B+Ours 0.7403 1.4136 0.3106 0.3411 0.5821
w/o Clustering 0.7271 1.3672 0.3094 0.3403 0.5821
4 Infinity2B+Ours 0.7437 1.4605 0.3112 0.3414 0.5821
w/o Clustering 0.7369 1.4323 0.3103 0.3409 0.5830
Table 14: Ablation study on clustering.
G.5 $\lambda$ Setting
In Table˜15, we exhibit the results using different temperature $\lambda$ in the resampling process with fixed clustering operations. Intuitively, higher temperature promotes the expression of intermediate states with higher potential scores and prevents superior samples. However, excessively high temperatures can also widen the gap between intermediate states with the highest scores and those with scores that are only slightly lower. This can directly inhibit the generation of these slightly lagging intermediate states, which may ultimately become the optimal results. As shown, though there is a steady increase in ImageReward, $\lambda=10.0$ falls behind $\lambda=5.0$ with $N=4$ in GenEval.
NN Lambda GenEval ImageReward HPS CLIP Aesthetic